Bwindi Impenetrable Forest (01.01.2016 to 03.31.2016)

You are here

PDF versionPDF version
Year: 
2016
Quarter: 
January - March
Communications activities: 

None

Unusual Events at the TEAM Site: 

None

New Species: 

None

Protocol Activities: 
Vegetation Plots
Re- measurement of the 5 plots that were not measured in the first quarter continued.  A total of 3091 trees were re-measured, tagged, mapped and identified to species level. 52 of the 3091 trees were new recruits. This brings the total to 3426 trees measured in this year.  The data for the plots has been entered in a spread sheet and will be uploaded in the TEAM portal in April 2016. 
 
Climate station
 Climate data for January to March 2016 has not been collected consistently and hence not been uploaded. This is because the battery power was fluctuating a lot and in order to avoid damage to the data logger, the climate station was turned off for sometime until the cause of the problem could be identified.  It was later found out that the solar panels were not providing enough energy, a new solar panel was installed and the battery is now functioning well.  Unfortunately, another problem with the temperature/humidity sensors arose. Two of the sensors had strange readings for sometime. After they were reconnected, one sensor ‘s readings normalised and the other continued to give wrong readings.  We tried all possible technical reconnections but without success. James from Conservation International thought that the problem could be with the calibration. Our next course of action will be to re-calibrate the sensors. 
 
Camera traps
There was no camera trapping activity. Camera trapping activity is scheduled for the beginning of May 2016. Eighteen new HC500 Hyper Fire Semi-Covert IR camera traps were purchased from RECONYX. These cameras will solve to a great extent camera trap failure challenges we have experienced in the past two years. 
 
Protocol Problems: 

 The battery power for the climate station was fluctuating  and droping a lot between January and March 2016. In order to avoid damage to the data logger, the climate station was turned off for sometime until the cause of the problem could be identified.  It was later noticed that the solar panels were not providing enough energy, a new solar panel was installed and the battery is now operating well.  Unfortunately, another problem with the temperature/humidity sensors arose. Two of the sensors provided wrong readings for sometime. After they were reconnected, one sensor ‘s readings normalised and the other continued to have strange readings.  We have made all possible reconnections but without success. James from Conservation International thinks that the problem could be with the calibration and has advised us to reset the calibration. Our next course of action will be re-calibration of the sensors. 

Schedule Problems: 

None

Logistical Problems: 

None

Suggestions: 

None